OpenStudy (jack1):

Hi all... have a bit of a challengeing question if anyone has some time to help please...? i'm totally lost. Find the inverse Laplace transform of the following function. \[\large F(s) = \frac {10}{s(s+2)(s+3)^2}\]

OpenStudy (jack1):

\[\large F(s) = \frac {10}{s(s+2)(s+3)^2}\]

OpenStudy (jack1):

@ganeshie8 or @skullpatrol ...if u guys have a spare couple of minutes, wood love a hand please?

terenzreignz (terenzreignz):

This looks nothing more than an incredibly tedious partial fraction decomposition problem :3

OpenStudy (jack1):

ok... but u lost me at decomposition

terenzreignz (terenzreignz):

partial fractions. Just split the bloody bugger into partial fractions. Crud... this is going to be tedious XD

terenzreignz (terenzreignz):

\[\Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2}\]

OpenStudy (jack1):

there's no way i can just split it somehow and use the laplace transform tables to reverse it somehow?

terenzreignz (terenzreignz):

This... IS splitting it somehow :D

terenzreignz (terenzreignz):

or something

terenzreignz (terenzreignz):

I haven't done Laplace transforms in about 18 months XD

OpenStudy (jack1):

\[\Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2}\] is this one too many? should the C/s+3 disappear?

terenzreignz (terenzreignz):

Let's try. Maybe I did overdo things XD

terenzreignz (terenzreignz):

It is rather daunting, isn't it? :D \[\Large \frac{A(s+2)(s+3)^2+ Bs(s+3)^2 + Cs(s+2)}{s(s+2)(s+3)^2}\]

terenzreignz (terenzreignz):

Wonder what this tells me... A+B = 0

OpenStudy (jack1):

whoa...whoa... that's ... ummm... huge

OpenStudy (jack1):

wait... why does A + B = 0 sorry man, i'm missing some logic thing there

terenzreignz (terenzreignz):

Because I said so. Isn't that logic enough? :)

terenzreignz (terenzreignz):

LOL kidding. Hang on...

OpenStudy (jack1):

lawl ;P

terenzreignz (terenzreignz):

When you evaluate all that out, the coefficient of \(x^3\) is going to be A+B. Try it.

terenzreignz (terenzreignz):

Convinced?

terenzreignz (terenzreignz):

Actually, I don't think this works at all. We need that D. Because this'll probably give us an overdefined system.

OpenStudy (jack1):

no, sorry, haven't dont partial fractions in ...11 months... where does x come in again?

terenzreignz (terenzreignz):

Oi... you better brush up on partial fractions, mate :P Laplace transforms crawl with them XD

terenzreignz (terenzreignz):

\[\Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2}\]

terenzreignz (terenzreignz):

Don't worry.. it's not THAT hard... hopefully XD \[\large F(s) = \frac{A(s+2)(s+3)^2 + Bs(s+3)^2 + Cs(s+2)(s+3) + Ds(s+2)}{s(s+2)(s+3)^2}\]

terenzreignz (terenzreignz):

crud... \[F(s) = \frac{A(s+2)(s+3)^2 + Bs(s+3)^2 + Cs(s+2)(s+3) + Ds(s+2)}{s(s+2)(s+3)^2}\]

OpenStudy (jack1):

hang on... u've lost me again... s(s+2)(s+3)(S+3)^2 doesnt equal s(s+2)(s+3)^2

terenzreignz (terenzreignz):

Don't be silly... true enough, if you multiply the denominators, it doesn't give you s(s+2)(s+3)^2 However, when adding fractions, it's not always the product of the denominators you get, but their LEAST COMMON DENOMINATOR. Which happens to be s(s+2)(s+3)^2 If you don't believe me, try separating \[F(s) = \frac{A(s+2)(s+3)^2 + Bs(s+3)^2 + Cs(s+2)(s+3) + Ds(s+2)}{s(s+2)(s+3)^2}\] And you'll definitely arrive at \[\Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2}\] after some cancellations.

OpenStudy (jack1):

ahhh. k with u now, cheers for clearing that up tj

terenzreignz (terenzreignz):

Okay, now to the hard part... I hope you're ready for this :D \[F(s) = \frac{A(s+2)(s+3)^2 + Bs(s+3)^2 + Cs(s+2)(s+3) + Ds(s+2)}{s(s+2)(s+3)^2}\]

terenzreignz (terenzreignz):

I can deduce A+B+C = 0 Any objections? :D

terenzreignz (terenzreignz):

But then again, my charge isn't really to solve this problem myself, than to ensure that you solve it. So... do you want to evaluate that long and scary polynomial? :D

OpenStudy (jack1):

some... just the deduction. A+B+C = 0 because...?

terenzreignz (terenzreignz):

Okay, good question. We need to find the coefficients of s^3, s^2, s, and the constant. So, to find the coefficient of s^3, say, in the first term we have (s+2)(s+3)^2 This'll be long and hard, but it'll inevitably start with an s^3 right? (s+2)(s+3)^2 = s^3 +...(some stuff) catch me so far?

OpenStudy (jack1):

yep, following.

OpenStudy (jack1):

s^3 + 8s^2 + 21s + 18

terenzreignz (terenzreignz):

So, the first term is A(s^3 + <some other stuff> ) Using similar logic, the second and third terms respectively become B(s^3 + <some other stuff> ) C(s^3 + <some other stuff> ) Don't worry about the other stuff just yet, for now, just concern yourself with the coefficient of s^3. So, we deduce (because in the fourth term, involving D, you don't get a s^3) that the coefficient of s^3 is A+B+C right?

OpenStudy (jack1):

ah, ok, understand the logic now... soory, u took some leaps before i didn't intuitively follow, with u now tj, cheers

terenzreignz (terenzreignz):

Now the harder bits... the coefficients of s^2 and s... This'll be a bloody nightmare XD We already know that in the first term (the A-bit) the coefficient of s^2 is 8 (as you already graciously figured :P )

terenzreignz (terenzreignz):

In the second bit, thankfully, it IS simpler. We have a square of a binomial, (s+3)^2 so we simply have to find the coefficient of the s in this binomial, which is 6. And that's the coefficient of s^2 in the second part (note the extra s outside)

OpenStudy (jack1):

\[s^3+6 s^2+9 s ~~is ~~B\]

terenzreignz (terenzreignz):

Well, I suppose I could leave the hard parts to you :D

terenzreignz (terenzreignz):

But I figured out it's 6 first >:) #OverlyCompetitive

terenzreignz (terenzreignz):

5 in the third.

OpenStudy (jack1):

aww... dont b like that... r u still hung up on the malaysian/phillipino confusion from earlier ? ;P

terenzreignz (terenzreignz):

Of course not. You know I respect you way too much to be worked up over a trivial error like that :D

OpenStudy (jack1):

sweeeet, and ditto dude

terenzreignz (terenzreignz):

Now, back to work. No objections to 5 being the coefficient of s^2 in the third term, I suppose?

OpenStudy (jack1):

nah, all good, its 5

terenzreignz (terenzreignz):

And what about the fourth? You didn't forget the fourth, did you? :)

OpenStudy (jack1):

1

OpenStudy (jack1):

1s^2 + 2s

OpenStudy (jack1):

winning ;D

terenzreignz (terenzreignz):

there we go. So we can deduce that 8A + 6B + 5C + D = 0 So far, so good?

OpenStudy (jack1):

yep, with u

terenzreignz (terenzreignz):

No you're not, you're like, miles away XD Anyway, now find the coefficients of the lone s. :P

OpenStudy (jack1):

21a, 9b, 2d and...

OpenStudy (jack1):

6c?

terenzreignz (terenzreignz):

Bingo :P 21A + 9B + 6C + 2D = 0 What about the remaining terms... the constants? :)

OpenStudy (jack1):

18,0,0,0.

terenzreignz (terenzreignz):

Yup. So we now have four unknowns, A,B,C,D and four equations. Now solve for A, B, C, and D.

terenzreignz (terenzreignz):

Hey, you were right, why delete that? XD

terenzreignz (terenzreignz):

18A = 10

OpenStudy (jack1):

shame and low self confidence...?

OpenStudy (jack1):

A = 5/9, B = -5, C = 40/9, D = 10/3 ...?

terenzreignz (terenzreignz):

that was fast :D

OpenStudy (jack1):

it was simultaneous eqn, used a website, dont need to relaern that.

terenzreignz (terenzreignz):

Okay... unfair. >:D Anyway, we now have values for A,B,C,D that should pretty much solve all your problems XD

OpenStudy (jack1):

wait... not really

OpenStudy (jack1):

if a + b + c = 0 and A + B + C + D = 10 then d = 10 but d = 10/3...?

OpenStudy (jack1):

no, nvmd, understood the glaring flaws in that logic, sorry

terenzreignz (terenzreignz):

I was going to say "Do I have to dignify that with a response" but decided against it :D

OpenStudy (jack1):

also glad u held back ;D

terenzreignz (terenzreignz):

Technically, I didn't, since you knew what I was going to say. Anyway, I take it you'll have no further problems in getting the inverse Laplace transform from here on in?

OpenStudy (jack1):

[\Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2}] [\Large F(s) = \frac{5/9}{s}+ \frac{-5}{s+2}+ \frac{40/9}{s+3}+ \frac{10/3}{(s+3)^2}] then pick and choose from the table whatever suits this?

OpenStudy (jack1):

[\Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2} \] [\Large F(s) = \frac{5/9}{s}+ \frac{-5}{s+2}+ \frac{40/9}{s+3}+ \frac{10/3}{(s+3)^2}\] then pick and choose from the table whatever suits this?...whay didn't that latexx properly...?

terenzreignz (terenzreignz):

I suggest you get the inverse Laplace first in terms of A,B,C, and D first, and then replace only AFTERWARDS. Less confusing, in my opinion.

terenzreignz (terenzreignz):

And you forgot the backslash \ in the beginning of those equations...

OpenStudy (jack1):

\[Large F(s) = \frac{A}{s}+ \frac{B}{s+2}+ \frac{C}{s+3}+ \frac{D }{(s+3)^2} \] \[Large F(s) = \frac{5/9}{s}+ \frac{-5}{s+2}+ \frac{40/9}{s+3}+ \frac{10/3}{(s+3)^2}\] then pick and choose from the table whatever suits this?... much better

terenzreignz (terenzreignz):

Well, I suppose you could work with this too. Yeah, you do that :D

OpenStudy (jack1):

cheers for all the help tj, really appreciate it hey!

terenzreignz (terenzreignz):

Of course you do >:)

OpenStudy (jack1):

5/9 ? \[\large -5e^{-2t}\] \[\Large \frac{40e^{-3t}}{9}\] \[\Large \frac{10e^{-3t}}{3} t\]

OpenStudy (jack1):

did i choose the right ones?

terenzreignz (terenzreignz):

Yup. I sense a little doubt with your first term, however... :P Is this what you call confidence? ;)

OpenStudy (jack1):

no, that's the shame and self loathing again... :'(

OpenStudy (jack1):

;P

OpenStudy (jack1):

so i just add them together into 1 equation from here then, yeah?

terenzreignz (terenzreignz):

one expression, and that's your answer, yes :)

OpenStudy (jack1):

back now.... sweeeeeeeeeet!